OpenStudy (anonymous):

find all the currents and the potential differences in the following circuit using Kirchhoff Rules

OpenStudy (anonymous):

i'll attach the paper soon :)

OpenStudy (anonymous):

@ajprincess @ghazi @shubhamsrg

OpenStudy (anonymous):

i have already tried but my equations are unsolvable :( should i attach my work too?

OpenStudy (vincent-lyon.fr):

Call i1 current in top loop and i2 current in bottom loop and write the 2 loop equations. Then we'll see what goes wrong.

OpenStudy (anonymous):

@Vincent-Lyon.Fr sorry am back :)

OpenStudy (anonymous):

ok so i keep the same direction?

OpenStudy (vincent-lyon.fr):

Keep the anti-clockwise direction for instance.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but we are going clockwise according to the arrow..

OpenStudy (anonymous):

this is what i did, i named A and B junction points.. A) I3+I2-I1 B)I1-I2-I3 is that correct so far?

OpenStudy (vincent-lyon.fr):

Yes, but this is just another way of writing i3 = i1 - i2 Write \(\Sigma\) (potential differences) = 0 is each loop.

OpenStudy (vincent-lyon.fr):

Sorry, go clockwise!

OpenStudy (anonymous):

ok then our loop equation then are the following: 1) U+I1R6-U2-I1R1-I3R3=0 5+3I1-3-5I3-4I1=0 2) I2(R2+R4+R5)+I3R3=0 9I2+5I3=0

OpenStudy (anonymous):

its ok, i'll just leave it and review other things :) thank you for the help though

OpenStudy (vincent-lyon.fr):

No, within a loop, the sign before any \(R_n\;i_{loop}\) will always be the same. Loop 1 must go: \(4i_1-5+3i_1+3+5(i_1-i_2)=0\)

OpenStudy (vincent-lyon.fr):

Loop 2 is: \(6i_2+1i_2+2i_2+5(i_2-i_1)=0\)